Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2001-237_2001-243

Mission Operations Days: 2001/237 to 2001/243
August 25, 2001 00:00:00 GMT to August 31, 2001 00:00:00 GMT


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
MODIS is in Science Mode - A Side
ComponentState(s)Comment
BlackbodyA On (290 K); B OffNominal
Calibration ElectronicsA On; B OffNominal
Control ProcessorA On; B OffNominal
Nadir DoorUnlatched, OpenNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA On; B OffNominal
FDDI PortA On; B OffNominal
FIFO Memory#1 & #2 On; #3 & #4 OffNominal
Format ProcessorA On; B OffNominal
Power Supplies#1 On; #2 OffNominal
PV VIS FPAA On; B OffNominal
PV NIR FPAA On; B OffNominal
PV SMIR FPAA On; B OffNominal
PV LWIR FPAA On; B OffNominal
PC LWIR FPAA On; B OffNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOn (83 K)Nominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA On; B OffNominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA Off; B OnNominal
Flight SoftwareRev BD + 2 set of patches
  • FR Error Counting and Statistics (2 patches)
  • FR Band 5 Gain Change (20 commands)
Nominal
Inhibit Ids SetNoneNominal
TMON 66 (SRCA A)DisabledNominal
TMON 67 (SRCA B)DisabledNominal
TMON 68 (FIFO)EnabledNominal
TMON 69 (EPS Current)DisabledNominal

This Week's Completed MODIS Activities:

Saturday, August 25, 2001

None

Sunday, August 26, 2001

None

Monday, August 27, 2001

None

Tuesday, August 28, 2001

one

Wednesday, August 29, 2001

2001/241 20:21:04 ATC OA-16: SD/SDSM Screened
2001/241 21:59:57 ATC OA-15: SD/SDSM Open

Thursday, August 30, 2001

None

Friday, August 31, 2001

None

This Week's Scheduled MODIS Activities Not Completed:

The formatter memory loads were not performed to clear the existing error addresses/counts.

Upcoming MODIS Events:

Friday, August 31, 2001

None

Saturday, September 1, 2001

None

Sunday, September 2, 2001

None

Monday, September 3, 2001

None

Tuesday, September 4, 2001

Anticipated day to perform memory load to clear FR error/count buffers

Wednesday, September 5, 2001

Two orbits with SMIR Itwk/Vdet of 79/110
SD/SDSM calibrations

Thursday, September 6, 2001

None

Friday, September 7, 2001

None

Maneuvers:

The next Lunar Calibration roll maneuver is scheduled for day 2001/250 (September 7, 2001) with data collection from 20:03:47 to 20:10:17. The roll angle will be -7.5106 degrees.

MODIS Anomalies:

The rate of formatter “error events” has increased during the past week to approximately 60-120 per day. There still have not been any noticeable impacts to operations.

General Instrument Comments:

MODIS is in Science Mode on the A-side with the SVD and NAD open, operating with the Formatter Resolution Flight Software Patches (FRFSP).

Note: SD/SDSM calibrations were planned to be reduced to once per two weeks. This plan has been revised; SD/SDSM calibrations will continue to be scheduled once each week until the end of November.

MODIS Telemetry Trends:

Trending has resumed with the instrument on the A-side. With the currently decreasing beta angle, the amount of focal plane heater margin is increasing. The intermediate stage cooler temperature is stable.

Non-MODIS Significant Events:

None.

Limited Life Item Status:

SRCA 10W Lamp #1: 189.1 of 500 hours
SRCA 10W Lamp #2: 144.8 of 500 hours
SRCA 10W Lamp #3: 154.0 of 500 hours
SRCA 10W Lamp #4: 61.5 of 500 hours

SRCA 1W Lamp #1: 558.0 of 4000 hours
SRCA 1W Lamp #2: 276.3 of 4000 hours

Solar Diffuser Door: 1691 of 3022 Movements
Nadir Aperture Door: 532 of 1316 Movements
Space View Door: 437 of 1316 Movements